OpenStudy (dumbsearch2):

It seems so crazy and IMPOSSIBLE to answer... Whoever will answer this correctly (I guarantee you, there IS an answer)... Is a GENIUS! A elongated pineapple is placed over the top of a prune, the same amount of elongated pineapple on either side of the prune. The elongated pineapple weighs one-third pound per foot. On one end of the elongated pineapple hangs Sir. Doofus the second holding a banana, and on the other end is a weight equal in weight to Sir. Doofus the second. The banana weighs two ounces per inch. The length of the elongated pineapple (in feet) is equal to the age of Sir. Doofus the second (in years), and the weight of Sir. Doofus the second (in ounces) is as much as the age of Sir. Doofus the second's girlfriend(s) (in years). The combined ages of Sir. Doofus the second and his girlfriend(s) are thirty years. The weight of the banana plus half of the weight of Sir. Doofus the second is one-fourth as much as the sum of the weights of the weight and the elongated pineapple, where all weights are in the same units. Sir. Doofus the second's girlfriend(s) is half as old as Sir. Doofus the second will be when he is three times as old as his girlfriend(s) when she was half as old as Sir. Doofus the second will be in when he is twice as old as he is now. How long is the banana?

OpenStudy (anonymous):

1 foot

OpenStudy (dumbsearch2):

Hint: the answer is not an integer.

OpenStudy (dumbsearch2):

Wrong @INeedSomeHelp!

OpenStudy (anonymous):

I don't get it. This makes no sense...

OpenStudy (dumbsearch2):

It does. It's just so super complex, I doubt anyone will get it here...

OpenStudy (anonymous):

Ya.

OpenStudy (jtvatsim):

lol, someone has too much time on their hands to come up with this question in the first place. :)

OpenStudy (timaashorty):

This is so confusing. it needs thinking, patience and time which I don't have right now lol @dumbsearch2 do you know the answer?

OpenStudy (dumbsearch2):

Yes, should I give it + an explanation? @timmashorty

OpenStudy (timaashorty):

Not yet, lets see if someone can answer it. I know I can't though. ;c Give it an hour. The unfinished/unclear sentences are the ones that confused me: Doofus the second's girlfriend(s) (in years). The weight of the banana plus half of the weight of Sir.

OpenStudy (timaashorty):

beautifulmelodies; you know it?

OpenStudy (solomonzelman):

I am too lazy to read :)

OpenStudy (timaashorty):

Dear 99's, @mathstudent55 @whpalmer4 @aaronq @Compassionate @beccaboo333 @timo86m @surjithayer Are you able to do this?

OpenStudy (kainui):

I can do this I won't though, on account of having a life.

OpenStudy (anonymous):

I am not reading all that.

OpenStudy (timaashorty):

these excuses, -.- D: if your not gonna do it, no comment.

OpenStudy (timaashorty):

@dumbsearch2 I want know the answer and explanation. :D

OpenStudy (timaashorty):

*to

OpenStudy (timaashorty):

@satellite73

OpenStudy (anonymous):

It's just a very long algebra problem.

OpenStudy (dumbsearch2):

let's let W = weight of Sir. Doofus the second = weight of the weight (in pounds) B = length of the banana (in inches) M = age of Sir. Doofus the second's girlfriend R = length of the elongated pineapple (in inches) T = age of Sir. Doofus the second P = weight of the banana (in pounds) H = weight of the elongated pineapple (in pounds) there's a lot of information, here, but it's hard to untangle. R/12 = T W*16 = M M+T = 30 P = 2B/16 H = (R/12)*(1/3) P + W/2 = 1/4(W + H) the last sentence is particularly hard to unravel. working backwards: "...when she was half as old as Sir. Doofus the second will be when he is twice is old as he is now" means: "when she is the same age as Sir. Doofus the second is now." so (the age) ..."Sir. Doofus the second will be when he is three times as old when..." means when Sir. Doofus the second is three times as old as his current age. so Sir. Doofus the second's girlfriend is half as old as Sir. Doofus the second will be when he is 3 times his current age: M = 1/2(3T) = 3T/2. so 3T/2 + T = 30 3T + 2T = 60 5T = 60 T = 12, Sir. Doofus the second is 12, his girlfriend is 18. this means the elongated pineapple is 12 feet long, or 144 inches long. this in turn, means the elongated pineapple weighs 4 pounds. now Sir. Doofus the second's girlfriend is 18, so the weight of Sir. Doofus the second is 18 ounces, or 9/8 pounds (this is also the weight of the weight). now P + W/2 = 1/4(W + H) and we already know W and H. P + (9/8)/2 = 1/4(9/8 + 4), so P = 9/32 + 1 - 9/16 = 9/32 + 32/32 - 18/32 = 23/32 pounds, or 11.5 ounces for the weight of the banana finally, the banana is 2 ounces per inch, so we have 1/2 inch per ounce (if 2B/16 = P, then B = (16P)/2, and 16P is the weight of the banana in ounces) thus the length of the banana is (1/2)(11.5) = 5.75 inches long

OpenStudy (dumbsearch2):

Excuse the long response... :-)

OpenStudy (timaashorty):

o m gee. that just needed deep thinking ;o But who's gonna end up eating that banana? I honestly thought it was going to be waaay bigger than that.

OpenStudy (anonymous):

best response me everyone

OpenStudy (timaashorty):

lol